# TableMind: An Autonomous Programmatic Agent for Tool-Augmented Table Reasoning

### v2

@@ Abstract.
Table reasoning is crucial for leveraging structured data in domains such as finance, healthcare, and scientific research. While large language models (LLMs) show promise in multi-step reasoning, purely text-based methods often struggle with the complex numerical computations and fine-grained operations inherently required in this task. Tool-integrated reasoning improves computational accuracy via explicit code execution, yet existing systems frequently rely on rigid patterns, supervised imitation, and lack true autonomous adaptability. In this paper, we present TableMind, an LLM-driven table reasoning agent that (i) autonomously performs multi-turn tool invocation, (ii) writes and executes data-analyzing code in a secure sandbox environment for data analysis and precise numerical reasoning, and (iii) exhibits high-level capabilities such as planning and self-reflection to adapt strategies. To realize these capabilities, we adopt a two-stage fine-tuning paradigm built on top of a powerful pre-trained language model: supervised fine-tuning on high-quality reasoning trajectories to establish effective tool usage patterns, followed by reinforcement fine-tuning to optimize multi-objective strategies. In particular, we propose Rank-Aware Policy Optimization (RAPO), which increases the update weight of high-quality trajectories when their output probabilities are lower than those of low-quality ones, thereby guiding the model more consistently toward better and more accurate answers. Extensive experiments on several mainstream benchmarks demonstrate that TableMind achieves superior performance compared to competitive baselines, yielding substantial gains in both reasoning accuracy and computational precision. To ensure the reproducibility of our results, we have made our experiment codes publicly available at https://github.com/lennendd/TableMind-Repo.
Table reasoning, Programmatic Agent, LLMs
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Tabular data (Hollmann et al., 2025; Ye et al., 2025) serves as a cornerstone for structured information across diverse domains, from scientific research(Lu et al., 2023) and clinical reporting(Pampari et al., 2018) to financial analysis(Chen et al., 2022). The combination of this data’s widespread use and recent breakthroughs in artificial intelligence (Lu et al., 2025; Cheng et al., 2025) has sparked significant academic and industrial interest in automated table reasoning(Wang et al., 2024c; Zhang et al., 2020). Effectively solving foundational tasks like table question answering(Cheng et al., 2021; Pasupat and Liang, 2015) or table fact verification(Aly et al., 2021; Chen et al., 2019) over tables presents a complex challenge, as it requires a combination of different skills: the precise retrieval of specific facts, robust natural language understanding, and accurate numerical computation.
The recent evolution of LLMs(Cheng et al., 2024; Zhai, 2024) has marked a paradigm shift in natural language processing. Techniques like Chain-of-Thought(Wei et al., 2022) prompting have unlock powerful multistep reasoning abilities. However, these models falter when confronted with quantitative tasks grounded in structured data, as their internal mechanisms for numerical computation and symbolic logic are inherently limited. To address this, research has increasingly focused on frameworks in which LLMs act as a reasoning engine to generate code for external execution. This synergy between the high-level reasoning of an LLM and the precise computation of a code environment has dramatically improved the accuracy of table reasoning. To structure this synergy, many state-of-the-art methods have adopted workflow-driven methodologies (Yao et al., 2023; Wang et al., 2024b), such as those following a plan-then-execute paradigm. These frameworks impose a structured control-flow on the problem-solving process, guiding the model through a pre-defined sequence of steps.
However, the rigidity of these predefined workflows presents three potential constraints, hindering the development of truly autonomous reasoning. Firstly, their reliance on fixed templates or analysis patterns restricts flexibility when encountering diverse or unseen table structures and query types. Secondly, models trained via supervised imitation to mimic these fixed reasoning trajectories are often restricted in their ability to explore novel or more efficient strategies autonomously. Thirdly, and most critically, these systems lack a mechanism for iterative self-correction; they tend to follow a linear execution path and struggle to adapt when faced with unexpected execution errors or intermediate results. Metaphorically speaking, these methods behave more like clerks following a fixed checklist—the workflow—rather than expert analysts who can improvise, learn from mistakes, and refine their approach on the fly. This distinction highlights the need to move from structured workflows to dynamic agents(Huang et al., 2024).
To bridge this gap, we propose TableMind, a novel programmatic agent for tool-augmented table reasoning that achieves autonomous and adaptive problem-solving. At its core, TableMind operates in an iterative plan-action-reflect loop. It first decomposes a complex problem into a sequence of executable steps. It then generates and runs Python code to perform each step. Finally, it critically evaluates the execution feedback—whether a successful output or an error message—to reassess its progress and dynamically adjust its subsequent actions. This iterative cycle empowers the agent to solve tasks with the flexibility and resilience of a human data analyst, enabling it to autonomously navigate complex reasoning paths and recover from errors. To train TableMind, we propose a two-stage optimization framework that enables the agent to evolve from imitation learning to strategic autonomy. The first SFT stage, injects an inductive bias for tool use, allowing the model to master essential syntax and common calling patterns. The second stage leverages Reinforcement Fine-tuning (RFT) to transition the agent from knowing how to use tools to understanding when, why, and how to use them optimally. To guide this process effectively, we design a fine-grained, multi-objective reward function and introduce RAPO, a novel algorithm that weights learning based on trajectory quality to improve exploration efficiency in complex strategy spaces. Through this framework, TableMind demonstrates two key advantages. First, by reflecting on execution outcomes, it can dynamically adjust its plan, achieving high flexibility. Second, its self-correction loop allows the agent to recover from intermediate failures, ensuring more reliable final outputs. Extensive experiments on multiple benchmark datasets show that TableMind achieves significant performance gains over existing baselines. Comprehensive analyses further highlight the framework’s effectiveness, adaptability, and ability to generate logical and coherent reasoning paths.To summarize, our main contributions can be listed as follows:
  • We propose TableMind, a novel agent framework for table reasoning that seamlessly integrates autonomous planning, precise tool execution, and reflective adaptation, enabling flexible and robust problem-solving.
  • We design a two-stage training paradigm that effectively combines SFT for foundational knowledge acquisition with RFT for advanced strategy optimization, further enhanced by a novel rank-aware policy optimization method.
  • We conduct extensive and rigorous experiments demonstrating that TableMind consistently achieves state-of-the-art performance, clearly showcasing its superior adaptability and advanced reasoning capabilities across diverse and challenging table reasoning tasks.
@@ 2. Related Work
Table reasoning has been a prominent research area for the last decade. The rise of LLMs and agents has recently led to major breakthroughs in core components like natural language understanding and numerical computation. This section will focus on research related to table reasoning and LLM agents.
@@ 2.1. Table Reasoning
By presenting data in a structured format, tables serve as a powerful means for organizing, storing, and analyzing information. Inspired by the success of Masked Language Modeling in BERT(Koroteev, 2021), involves masking parts of the table. For instance, TaPas(Herzig et al., 2020) adopts this by requiring the model to reconstruct masked cells during pre-training. Pasta(Gu et al., 2022) and TUTA(Wang et al., 2021) extend this concept by proposing to mask entire columns or segments of the table. In contrast, TAPEX(Liu et al., 2021) pursues a different paradigm. It pre-trains an encoder-decoder model on a large, synthetic SQL dataset. This enables the model to function as a SQL executor, thereby achieving a deeper understanding of the tabular structure. With the emergence of LLMs, the focus shifted toward general-purpose systems that can perform multiple table reasoning tasks without task-specific retraining. Models such as TableLLaMA(Zhang et al., 2023), TableLLM(Zhang et al., 2024), and the TableGPT series(Zha et al., 2023; Su et al., 2024) adapt LLMs to structured data by combining language reasoning with table parsing, often achieving strong zero-shot and few-shot performance. Another line of work emphasizes workflow-driven methods. Chain-of-Table(Wang et al., 2024c) follows a data-flow paradigm, updating the table after each reasoning step, while PoTable(Mao et al., 2024) adopts a plan-then-execute strategy, decoupling planning from execution through external tools. Recent reinforcement learning frameworks have notably enhanced the reasoning capabilities of LLMs, as seen in DeepSeek-R1(Guo et al., 2025). Extending this line, Table-R1(Yang et al., 2025) applies rule-based reinforcement learning to structured tables. Overall, language models have become central to table reasoning(Wang et al., 2024a), underscoring their growing research significance.
@@ 2.2. LLM-based Agent
Large Language Models often struggle with complex reasoning or multi-step planning tasks when relying solely on purely text-based inference. To address these challenges, recent work has increasingly integrated external tool use and explicit planning strategies to augment model capabilities(Wang et al., 2024b). For example, some methods adopt an iterative data-flow approach that continuously updates an internal state after each reasoning step, while others follow a plan-and-execute paradigm—first generating a detailed plan and then invoking external tools to carry it out. However, such rigid, predefined pipelines can significantly limit flexibility, motivating the development of more dynamic agent frameworks that plan and select tools on the fly. One such approach is ReAct (Yao et al., 2023), which interleaves reasoning and action: the agent alternates between thinking and taking an action based on its current reasoning state, enabling continuous interaction with an external environment. Another example is CRITIC (Gou et al., 2023), where the agent learns via self-reflection by using external resources (e.g., search engines or knowledge bases) to validate and refine its decisions, effectively self-correcting mistakes. Building on these ideas, our work trains an LLM-based agent on a smaller-scale model to explore autonomous tool use. Despite progress in LLM agent research, the design of fully autonomous execution pipelines—such as enabling the agent to write and run code within a secure sandbox—remains largely unexplored.
@@ 3.2. Design Principles
Large language models demonstrate strong capabilities in understanding natural language, making them valuable tools for the complex task of table reasoning. This task requires interpreting structured data, executing precise computations, and synthesizing intermediate results, often necessitating interaction with external tools to ensure accuracy. However, current research presents a distinct dilemma, forcing a choice between two insufficient paradigms. On one hand, many existing approaches omit tool usage entirely, treating reasoning as a single-pass generation process. These methods lack mechanisms for external verification or self-correction, often failing on tasks that require complex computation. On the other hand, the few methods that do incorporate tool interaction are heavily reliant on massive, proprietary models. Their tool invocation patterns are often rigid and pre-defined, lacking the capacity for dynamic reflection, and their dependence on large-scale models creates significant barriers to computational cost and adaptability. This landscape reveals a critical and unfilled research gap: the absence of systems that combine autonomous tool invocation and reflection with lightweight, accessible models.
Figure 2. The overall training pipeline for TableMind. The process begins with Prompt Building and SFT to warm up the model by providing it with a strong initial policy. Subsequently, RFT with RAPO is applied to significantly enhance its generalization capability. The final model is then deployed for Inference and evaluated for accuracy.
@@ 4. TableMind
This section describes the construction of the TableMind, which is trained using a two-stage process of Supervised Fine-tuning and Reinforcement Fine-tuning. This allows it to implement a plan-action-reflect loop for flexible and autonomous table reasoning.
@@ 4.1. Overview
TableMind consists of a two-stage training framework for table reasoning with LLMs. This framework leverages a structured prompt template to unify various table inputs and to teach the model how to understand task instructions. The first SFT stage aims to equip the model with the foundational ability to invoke tools. We use examples that distill from expert models as training data. The second stage employs RFT to guide the model in autonomously exploring superior tool-use strategies and more complex reasoning paths. We designed a fine-grained reward function tailored for table reasoning and introduced the RAPO method. By ranking and weighting reasoning trajectories of varying quality, RAPO efficiently optimizes decision sequences to strike a balance between final answer accuracy. An overview of the framework is provided in Figure  2.
4.2. Prompt Template
We design a structured prompt template to ensure consistent and accurate reasoning from the Large Language Model. Our prompt is constructed by combining two main components: a general Instruction block and a problem-specific table task block. The Instruction component defines the model’s persona and core operational methodology. It assigns the role of an expert AI data analyst and mandates a specific reasoning process: analyzing the problem, planning a strategy, executing code for data manipulation and reflecting the output. The table task block component provides the concrete details for each specific problem instance. It consists of the table content and the corresponding table question.
@@ 4.3. Supervised Fine-tuning
Applying reinforcement learning directly to complex tabular reasoning poses significant challenges—specifically, an overly broad search space and slow convergence. To address these issues, training begins with an essential SFT stage. This stage serves as a warm-up phase, aligning the model’s output format, enabling it to master the basic syntax of the plan–action–reflect cycle, and establishing an initial, reasonably stable policy for generating valid and logically coherent code. This design alleviates the instability associated with starting from a cold policy that lacks basic tool-use capabilities. We construct a high-quality dataset of reasoning trajectories for SFT using an iterative, multi-turn framework. At its core, this framework employs a knowledge distillation(Hinton et al., 2015) approach, aiming to convert the programmatic knowledge and reasoning processes of an expert model into structured training samples. First, we prompt the expert model to generate a plan and a corresponding initial executable action-a code snippet. This action is then executed in a local environment, and the resulting output or error message is returned to the model as an observation. Upon receiving this feedback, the model enters a reflection phase to evaluate the outcome of the preceding action, based on which it refines its plan and generates the next action. This interactive loop of plan-action-reflect continues until the task is completed and a final answer is obtained. All collected trajectories undergo a rigorous calibration process. We compare the final output of each trajectory against a predefined ground-truth answer, retaining only those that accurately lead to the correct solution. This filtering ensures that the final dataset comprises exclusively high-quality, validated reasoning paths. After collecting the training data, we fine-tune the model using SFT with a cross-entropy loss function. The initial SFT phase serves as a crucial preparatory step, lowering the overhead for the subsequent RFT. While this stage successfully imparts foundational instruction-following skills and nascent reasoning abilities, the reasoning itself is not yet robust. It largely reflects a form of pattern memorization from the training data, exhibiting poor generalization to out-of-distribution tasks. Therefore, this SFT-induced reasoning can be characterized as a form of memory that lacks the ability to generalize, necessitating further refinement through RFT.

### v3

@@ Abstract.
Table reasoning requires models to jointly perform comprehensive semantic understanding and precise numerical operations. Although recent large language model (LLM)-based methods have achieved promising results, most of them still rely on a single-turn reasoning paradigm that processes flattened tables in a single forward pass. This paradigm suffers from inherent limitations, including context overflow on large tables, weak sensitivity to continuous numerical values, and the absence of explicit tool-use and reflection. In this paper, we propose TableMind, a tuning-based autonomous programmatic table agent that simulates the human-like cognitive schema of the multi-turn interaction within a lightweight LLM. Instead of adopting a training-free workflow design, TableMind learns to internalize planning, action, and reflection through a principled two-stage training strategy. To bootstrap structured table reasoning capabilities, we construct and filter high-quality reasoning data for the supervised fine-tuning (SFT) stage. To enable precise code generation, we introduce a designed multi-perspective reward scheme and a novel optimization objective in the reinforcement learning (RL) stage. Extensive experiments on diverse benchmarks demonstrate that TableMind consistently outperforms previous baselines, validating the effectiveness of training autonomous agents to improve overall performance.  [^1] Our codes are available at https://github.com/lennendd/TableMind.
Table Reasoning, Autonomous Agent, Programming
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USA.booktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAisbn: 979-8-4007-2292-9/2026/02doi: 10.1145/3773966.3777932ccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Tables are one of the most fundamental forms of structured data, widely used across scientific research  (hollmann2025accurate), healthcare management (pampari2018emrqa), and business intelligence (chen2022convfinqa). A large amount of critical knowledge is organized in tabular formats, making table understanding and reasoning a practically important research problem (ye2025closer). Unlike unstructured text, table reasoning requires models to jointly handle comprehensive semantic understanding and precise numerical operations.
Early studies on table reasoning primarily rely on deep learning models, which focus on task-specific representation learning for structured tables (herzig2020tapas; hollmanntabpfn). These approaches model table structures and numerical relationships through masked pre-training or end-to-end learning paradigms (liu2021tapex). While these methods achieve strong performance on specific tasks, they often depend on extensive labeled supervision and generalize poorly to unseen table structures. The recent advances in large language models (LLMs) (cheng2024towards; zhai2024large) introduce new opportunities for table reasoning. By leveraging rich world knowledge and strong semantic reasoning capabilities, LLM-based approaches mainly reformulate table reasoning as a language-driven generative process, enabling unified reasoning and zero-shot generalization  (zhang2024tablellm).
Despite their effectiveness, most existing methods still adopt a single-turn reasoning paradigm, processing tables in a single forward pass, which leads to several fundamental limitations (cheng2025survey). First, many LLM-based approaches (tao2025values) flatten the entire table into textual input and directly feed it to the language model. This design is vulnerable to context overflow for large tables, leading the model to reason over incomplete information, particularly in wide or long tables. Second, LLM-based methods (zhou2025benchmarking) are inherently insensitive to continuous numerical values. While they excel at semantic reasoning, they often struggle with precise numerical operations such as arithmetic, comparison, and aggregation. Treating values as plain tokens hinders reliable numerical reasoning, leading to unstable outputs and critical calculation errors. Third, existing LLM-based methods (cheng2025agent) typically lack explicit tool-use and reflection. Most existing LLM-based methods mainly depend on the intrinsic capabilities of LLMs and lack explicit mechanisms for code generation, execution, and reflection. Overall, these limitations reveal a mismatch between single-turn LLM reasoning and the multi-turn interaction process required for reliable table reasoning.
These analyses motivate a closer examination of how human cognitively perform table reasoning, as illustrated in the left part of Figure 1. Given a query, humans typically adopt a systematic multi-stage process to derive answers from complex tables. Specifically, they analyze the query and decompose it into sub-tasks, locating relevant rows and columns, performing explicit numerical operations, checking intermediate results, and finally deriving a validated answer. A key characteristic of this process is that reasoning proceeds via explicit actions, such as calculations or numerical operations, while intermediate results are repeatedly verified (chen2019tabfact). This observation suggests that reliable table reasoning requires explicit planning, action, and reflection, rather than a single-turn generation over structured tables.
Motivated by the human-like cognitive schema of the multi-turn interaction, a direction is to leverage LLM-driven agents (huang2024understanding) to explicitly model such processes, as shown in the right part of Figure 1. In practice, LLMs are coordinated through predefined pipelines to perform such processes, with the most straightforward implementation being training-free workflow-based agents (yao2023react; wang2024survey). While effective for numerical reasoning, it comes with notable limitations (cheng2024towards; zhai2024large). Workflow-based agents typically rely on multiple LLM calls and external execution environments, resulting in computational overhead. Moreover, the need to expose tabular data to external models raises practical concerns regarding data privacy and security, especially in sensitive real-world applications. Given these limitations, we aim for a more meaningful and challenging goal: instead of relying on heavy workflows, we seek to train a lightweight table agent that intrinsically acquires human-like reasoning capabilities. Such an agent is expected to internalize planning, action, and reflection capabilities within an LLM, enabling efficient, privacy-preserving, and accurate table reasoning.
To achieve the above goal, we propose TableMind, a lightweight autonomous agent trained to internalize human-like cognitive schema of the multi-turn interaction. TableMind is trained using a principled two-stage strategy. In the first stage, we perform supervised fine-tuning with carefully constructed high-quality data, allowing the model to learn the basic capability of table reasoning, including problem decomposition, code generation, and numerical operations. To ensure the effectiveness of this stage, we adopt a strict data filtering strategy to extract reasoning trajectories from a teacher model. In the second stage, we further introduce a reinforcement learning strategy that encourages trial-and-error exploration. We design multi-view rewards together with a novel optimization objective to promote accurate, verifiable, and reflective reasoning trajectories. Through this two-stage training process, TableMind is able to perform multi-turn interaction reasoning, including planning, automatic program generation, numerical reasoning, and reflection in a lightweight manner. Extensive experiments on diverse public benchmarks demonstrate that TableMind consistently outperforms previous baselines, validating the effectiveness of training a lightweight table agent for accurate table reasoning.
Our main contributions are summarized as follows:
  • We propose TableMind, a lightweight autonomous agent that explicitly performs multi-turn interactive reasoning through planning, action, and reflection.
  • We introduce a two-stage training strategy combining supervised fine-tuning and reinforcement learning to simulate human-like table reasoning capabilities.
  • We conduct extensive experiments on multiple benchmarks, showing consistent improvements over previous methods.
@@ 2. Related Work
With recent advances in LLMs and agents, this section focuses on research related to table reasoning and LLM agents.
@@ 2.1. Table Reasoning
By presenting data in a structured format, tables serve as a powerful means for organizing, storing, and analyzing information. Inspired by the success of masked language modeling in BERT (koroteev2021bert), involves masking parts of the table. For instance, TaPas (herzig2020tapas) adopts this by requiring the model to reconstruct masked cells during pre-training. Pasta (gu2022pasta) and TUTA (wang2021tuta) extend this concept by proposing to mask entire columns or segments of the table. In contrast, TAPEX (liu2021tapex) pursues a different paradigm. It pre-trains an encoder-decoder model on a large, synthetic SQL dataset. This enables the model to function as a SQL executor, thereby achieving a deeper understanding of the tabular structure. With the emergence of LLMs, the focus shifted toward general-purpose systems that can perform multiple table reasoning tasks without task-specific retraining. Models such as TableLLaMA (zhang2023tablellama), TableLLM (zhang2024tablellm), and the TableGPT series (zha2023tablegpt; su2024tablegpt2) adapt LLMs to structured data by combining language reasoning with table parsing, often achieving strong zero-shot and few-shot performance. Another line of work emphasizes workflow-driven methods. Chain-of-Table (wang2024chain) follows a data-flow paradigm, updating the table after each reasoning step, while PoTable (mao2024potable) adopts a plan-then-execute strategy, decoupling planning from execution through external tools. Recent reinforcement learning frameworks have notably enhanced the reasoning capabilities of LLMs, as seen in DeepSeek-R1 (guo2025deepseek; luo2025time; wang2025can). Extending this line, Table-R1 (yang2025table) applies rule-based reinforcement learning to structured tables. Overall, language models have become central to table reasoning (wang2024tabletime), underscoring their research significance.
@@ 2.2. LLM-based Agent
LLMs often struggle with complex reasoning or multi-step planning tasks when relying on purely text-based inference. To address these challenges, recent work has increasingly integrated external tool use and explicit planning strategies to augment model capabilities (wang2024survey). For example, some methods adopt an iterative data-flow approach that continuously updates an internal state after each reasoning step, first generating a detailed plan and then invoking external tools to carry it out. However, such rigid, predefined pipelines can significantly limit flexibility, motivating the development of more dynamic agent frameworks that plan and select tools on the fly. One such approach is ReAct  (yao2023react), which interleaves reasoning and action: the agent alternates between thinking and taking an action based on its current reasoning state, enabling continuous interaction with an external environment. Another example is CRITIC (gou2023critic), where the agent learns via self-reflection by using external resources (e.g., search engines or knowledge bases) to validate and refine its decisions, effectively self-correcting mistakes. Building on these ideas, our work trains an LLM-based agent on a smaller-scale model to explore autonomous tool use. Despite progress in LLM agent research, the design of fully autonomous execution pipelines—such as enabling the agent to write and run code within a secure sandbox—remains largely unexplored.
Figure 2. Overview of the multi-stage training pipeline for TableMind. The process begins with Prompt Design and SFT to warm up the model by providing it with a strong initial policy. Subsequently, RFT with RAPO is applied to significantly enhance its generalization capability. The final model is then deployed for Inference and evaluated for accuracy.
@@ 3.2. Design Principles
LLMs demonstrate strong capabilities in understanding natural language, making them valuable tools for the complex task of table reasoning (zhang2020summarizing; zhang2025star). This task requires interpreting structured data, executing precise computations, and synthesizing intermediate results, often necessitating interaction with external tools to ensure accuracy. However, current research presents a distinct dilemma, forcing a choice between two insufficient paradigms. These methods lack mechanisms for external verification or self-correction, often failing on tasks that require complex computation. On the other hand, the few methods that do incorporate tool interaction (masterman2024landscape) are heavily reliant on massive, proprietary models. Their tool invocation patterns are often rigid and pre-defined, lacking the capacity for dynamic reflection, and their dependence on large-scale models creates significant barriers to computational cost and adaptability. This landscape reveals a critical and unfilled research gap: the absence of systems that combine autonomous tool invocation and reflection with lightweight, accessible models.
@@ 4. TableMind
This section presents the training pipeline of TableMind. It begins with the prompt template design, followed by the two-stage training strategy, and concludes with inference and evaluation.
@@ 4.1. Overview
An overview of TableMind is shown in Figure 2. We first design a Prompt Template to standardize diverse table inputs and task descriptions, guiding the model to reason by combining thinking and tool usage. TableMind uses a two-stage training strategy. The supervised fine-tuning (SFT) stage equips the model with basic tool-invocation and reasoning capabilities using trajectories distilled from teacher model. The reinforcement fine-tuning (RFT) stage allows the model to autonomously explore and refine reasoning strategies. To improve stability in reasoning tasks, we introduce a rank-aware policy optimization (RAPO) algorithm and a multi-perspective reward function, aligning model confidence with trajectory quality for stable policy improvement. During Inference and Evaluation, the model executes iterative reasoning over tables, generating actions, and deriving final answers.
4.2. Prompt Template Design
We propose a structured prompt template to ensure consistent and reliable reasoning for table-based tasks, aiming to reduce ambiguity in table understanding and to provide a stable interface for reasoning and tool usage. The prompt integrates high-level reasoning guidance with task-specific information to address the inherent complexity of table reasoning, which requires coordinated data interpretation, problem decomposition, and tool invocation. Concretely, it consists of a general Instruction block that defines the model’s role and enforces a fixed reasoning workflow, and a Table Task block that provides the table content and corresponding query. This design yields more stable reasoning trajectories and more accurate tool usage across diverse table reasoning tasks.
@@ 4.3. Supervised Fine-tuning
The SFT stage primarily warms up the agent’s interaction behavior by distilling high-quality trajectories and aligning the model with a structured reasoning paradigm.
4.3.1. Data Collection
We construct a high-quality dataset of reasoning trajectories for SFT using an iterative, multi-turn framework. At its core, this framework employs a knowledge distillation (hinton2015distilling) approach to transfer the programmatic knowledge and reasoning processes of an expert model into structured training samples. Specifically, we prompt the expert model to generate a plan and an initial executable action, which is then executed in a local environment. The resulting output or error message is returned to the model as an observation, enabling it to reflect on the outcome, refine its plan, and generate the next action. This interactive loop continues until the task is completed and a final answer is derived. All collected trajectories undergo a rigorous calibration process: we compare the final output of each trajectory against a predefined ground-truth answer and retain only those that lead to correct solutions. This filtering ensures that the resulting dataset consists exclusively of high-quality, validated reasoning paths.
4.3.2. Warm-up Training Stage
The SFT stage serves as a warm-up phase that aligns the model’s output format, enables it to master the basic syntax of the planning–action–reflection loop, and establishes an initial policy for generating valid and logically coherent code. This initial training phase acts as a preparatory step, substantially lowering the optimization overhead for subsequent RFT. However, while SFT successfully imparts foundational instruction-following capabilities and reasoning skills, the resulting reasoning behavior is not yet consistent. It primarily reflects pattern memorization from the training data and exhibits limited generalization to out-of-distribution tasks. Consequently, the reasoning ability induced by SFT can be viewed as a form of non-generalizable memory, necessitating further refinement through RFT.